# Euclid Quick Data Release (Q1)

### v1

@@ Appendix A Description of the spectroscopically identified bright quasar catalogue
The format of the catalogue of spectroscopically identified bright quasars from Euclid Q1 is shown in Table 5.
@@
Table 5: Format of the catalogue of spectroscopically identified bright quasars from Euclid Q1.777The catalogue described by Table 5 is available in its entirety in the electronic edition.
@@
7 sum_mask int Sum of the mask values of the spectrum within [12047, 18734]
8 n_invalid int Number of invalid pixels in the spectrum within [12047, 18734]
9 med_snr double Median signal-to-noise ratio of the spectrum within [12047, 18734]
@@
12 gaia_id long The associated GAIA source id
13 f_agn double AGN fraction of the source from VIS image (Q1-SP015)
@@
Appendix B Composite spectra generation with 1D drizzle
B.1 Details of the 1D drizzle method
We construct Euclid quasar composite spectra on a common rest-frame grid with a constant bin size of $\Delta\lambda=4\,\AA$ . The mapping preserves the integrated flux density in each bin through a linear, flux-conserving rebinning that partitions every input pixel across the overlapping output bins in proportion to the fractional overlap in wavelength. This procedure is mathematically equivalent to a one-dimensional version of the Drizzle algorithm (2002PASP..114..144F) and matches the smallest native rest-frame pixel size at the blue end of our spectra, while mildly oversampling the data at longer wavelengths.
Preprocessing and normalisation. Each spectrum is shifted to the rest frame using the adopted redshift, then converted to rest-frame flux density in $F_{\lambda}$ via $F_{\lambda,\mathrm{rest}}=(1+z)\,F_{\lambda,\mathrm{obs}}$. The spectra are rebinned onto the common rest-frame grid with the flux-conserving drizzle scheme described below. Each rebinned spectrum is then normalised to the running composite within the spectral overlap so that relative shapes are combined consistently; the same normalisation factor is applied to the spectrum’s per-pixel variances. We ignore pixels with non-finite or non-positive flux or variance values.
1D drizzle rebinning and variance propagation. Let $f_{p}$ and $\sigma_{p}^{2}$ be the flux and variance in input pixel $p$, whose wavelength interval is $\lambda_{p-1/2},\lambda_{p+1/2}]$ with width $\Delta\lambda_{p}=\lambda_{p+1/2}-\lambda_{p-1/2}$. The $j$-th output bin covers $\lambda_{j-1/2},\lambda_{j+1/2}]$ with width $\Delta\lambda_{j}$. We define the fractional overlap
@@ B.1 Details of the 1D drizzle method
Equal-weight mean and its uncertainty. For the mean composite we combine, in each bin $j$, all rebinned spectra that have valid data in that bin with equal weights. If $n_{j}$ spectra contribute and their rebinned fluxes and variances are $\hat{f}_{ij}$ and $\widehat{\sigma}_{ij}^{2}$, the mean composite and its formal uncertainty are
@@
where MAD is the median absolute deviation (LEYS2013764). For a univariate data set $X_{1}$, $X_{2}$, …, $X_{n}$, the MAD is defined as the median of the absolute deviations from the data’s median
@@
Median composite. We form a median composite by taking, in each bin, the median of the contributing $\hat{f}_{ij}$ values. We apply the same sigma clipping as done when computing the mean composite. We do not attach formal uncertainties to the median composite.
Geometric-mean composite for continuum shape. To preserve the average continuum slope of quasars, we also compute a geometric-mean composite on the same rebinned spectra and wavelength grid, again using equal weights. For contributors with positive flux,
@@
We mask bins where fewer than three spectra contribute or where negative flux values prevent a robust logarithmic mean.
Effect of wavelength bin size. In Sect. 4.3 we adopted a bin size of $\Delta\lambda=4\,\AA$ for constructing the Euclid mean composite spectrum. To assess the impact of this choice, we recompute the mean composite using coarser grids with $\Delta\lambda=8\,\AA$ and $13.4\,\AA$ , keeping all other steps of the stacking procedure fixed. As shown in Fig. 13, the continuum shapes of three mean composites are virtually identical in both the optical and near-infrared ranges, while the two larger bin sizes produce undersampled, flattened emission line peaks. This test demonstrates that the bin size of $\Delta\lambda=4\,\AA$ is favourable for the preservation of details of the composite spectra.
@@
B.2 Construction of the piecewise quasar template
@@
    2001AJ....122..549V for $\lambda_{\rm rest}<320$ nm;
@@
    Glikman_2006 for $\lambda_{\rm rest}>1550$ nm.
All three spectra are resampled onto a common rest-frame grid from 90 to 2100 nm with a spacing of 0.2 nm using a flux-conserving resampler (specutils.FluxConservingResampler). Before splicing, we scale the adjoining segments using the mean flux in 5 nm windows centred on each join. At 320 nm we scale our composite to match 2001AJ....122..549V; at 1550 nm we scale Glikman_2006 to match our composite. If $\langle F_{\mathrm{A}}\rangle$ and $\langle F_{\mathrm{B}}\rangle$ are the mean fluxes in the overlap window, the scale factor applied to $F_{\mathrm{B}}$ is $s=\langle F_{\mathrm{A}}\rangle/\langle F_{\mathrm{B}}\rangle$. After scaling, we combine the spectra piecewise at the join wavelengths (no additional cross-fade). The resulting template covers 90 to 2100 nm on a uniform grid.
Appendix C Example cutouts of quasars
@@
Figure 14: Euclid 1D spectra and imaging cutouts for a random sample of 10 low-redshift ($z<0.5$) sources. Each panel displays the 1D spectrum (top) and five imaging cutouts with $10^{\prime\prime}$ sizes (bottom) in the $I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ bands, as well as a VIS-$Y_{\scriptscriptstyle\rm E}$ composite. The composite image is generated by mapping the VIS and $Y_{\scriptscriptstyle\rm E}$ fluxes into the blue and red channels, respectively, with their mean used for green, and the VIS band used to define overall luminosity in the L*a*b* colour space (iso11664-4) to enhance morphological detail. Major emission lines detected in the wavelength range [12 047, 18 734] Å are marked.

### v2

@@ Appendix B Description of the spectroscopically identified bright quasar catalogue
The format of the catalogue of spectroscopically identified bright quasars from Euclid Q1 is shown in Table 5. The full catalogue will be available at https://cdsarc.cds.unistra.fr/.
@@
Table 5: Format of the catalogue of spectroscopically identified bright quasars from Euclid Q1.999The catalogue described by Table 5 is available in its entirety at the CDS.
@@
7 sum_mask int Sum of the mask values of the spectrum within [12047, 18734] Å
8 n_invalid int Number of invalid pixels in the spectrum within [12047, 18734] Å
9 med_snr double Median signal-to-noise ratio of the spectrum within [12047, 18734] Å
@@
12 gaia_id long The associated Gaia DR3 source id
13 f_psf double PSF fraction of the source from VIS image (Euclid Collaboration: Margalef-Bentabol et al. 2025)
@@
Appendix C Composite spectra generation with 1D drizzle
C.1 Details of the 1D drizzle method
We construct Euclid quasar composite spectra on a common rest-frame grid with a constant bin size of $\Delta\lambda=4\,\AA$ . The mapping preserves the integrated flux density in each bin through a linear, flux-conserving rebinning that partitions every input pixel across the overlapping output bins in proportion to the fractional overlap in wavelength. This procedure is mathematically equivalent to a one-dimensional version of the Drizzle algorithm (Fruchter and Hook 2002) and matches the smallest native rest-frame pixel size at the blue end of our spectra, while mildly oversampling the data at longer wavelengths.
Preprocessing and normalisation.
Each spectrum is shifted to the rest frame using the adopted redshift, then converted to rest-frame flux density in $F_{\lambda}$ via $F_{\lambda,\mathrm{rest}}=(1+z)\,F_{\lambda,\mathrm{obs}}$. The spectra are rebinned onto the common rest-frame grid with the flux-conserving drizzle scheme described below. Each rebinned spectrum is then normalised to the running composite within the spectral overlap so that relative shapes are combined consistently; the same normalisation factor is applied to the spectrum’s per-pixel variances. We ignore pixels with non-finite or non-positive flux or variance values.
The 1D drizzle rebinning and variance propagation.
Let $f_{p}$ and $\sigma_{p}^{2}$ be the flux and variance in input pixel $p$, whose wavelength interval is $\lambda_{p-1/2},\lambda_{p+1/2}]$ with width $\Delta\lambda_{p}=\lambda_{p+1/2}-\lambda_{p-1/2}$. The $j$-th output bin covers $\lambda_{j-1/2},\lambda_{j+1/2}]$ with width $\Delta\lambda_{j}$. We define the fractional overlap
@@ C.1 Details of the 1D drizzle method
Equal-weight mean and its uncertainty.
For the mean composite we combine, in each bin $j$, all rebinned spectra that have valid data in that bin with equal weights. If $n_{j}$ spectra contribute and their rebinned fluxes and variances are $\hat{f}_{ij}$ and $\widehat{\sigma}_{ij}^{2}$, the mean composite and its formal uncertainty are
@@
where MAD is the median absolute deviation (Leys et al. 2013). For a univariate dataset $X_{1}$, $X_{2}$, …, $X_{n}$, the MAD is defined as the median of the absolute deviations from the data’s median
@@
Median composite.
We formed a median composite by taking, in each bin, the median of the contributing $\hat{f}_{ij}$ values. We applied the same sigma clipping as done when computing the mean composite. We did not attach formal uncertainties to the median composite.
Geometric-mean composite for continuum shape.
To preserve the average continuum slope of quasars, we also computed a geometric-mean composite on the same rebinned spectra and wavelength grid, again using equal weights. For contributors with positive flux,
@@
We masked bins where fewer than three spectra contribute or where negative flux values prevent a robust logarithmic mean from being determined.
Effect of wavelength bin size.
In Sect. 4.3 we adopted a bin size of $\Delta\lambda=4\,\AA$ for constructing the Euclid mean composite spectrum. To assess the impact of this choice, we recomputed the mean composite using coarser grids with $\Delta\lambda=8\,\AA$ and $13.4\,\AA$ , keeping all other steps of the stacking procedure fixed. As shown in Fig. 14, the continuum shapes of three mean composites are virtually identical in both the optical and near-infrared ranges, while the two larger bin sizes produce undersampled, flattened emission line peaks. This test demonstrates that the bin size of $\Delta\lambda=4\,\AA$ is favourable for the preservation of details of the composite spectra.
@@
C.2 Construction of the piecewise quasar template
@@
    Vanden Berk et al. (2001) for $\lambda_{\rm rest}<320$ nm;
@@
    Glikman et al. (2006) for $\lambda_{\rm rest}>1550$ nm.
All three spectra are resampled onto a common rest-frame grid from 90 to 2100 nm with a spacing of 0.2 nm using a flux-conserving resampler (specutils.FluxConservingResampler). Before splicing, we scale the adjoining segments using the mean flux in 5 nm windows centred on each join. At 320 nm we scale our composite to match Vanden Berk et al. (2001); at 1550 nm we scale Glikman et al. (2006) to match our composite. If $\langle F_{\mathrm{A}}\rangle$ and $\langle F_{\mathrm{B}}\rangle$ are the mean fluxes in the overlap window, the scale factor applied to $F_{\mathrm{B}}$ is $s=\langle F_{\mathrm{A}}\rangle/\langle F_{\mathrm{B}}\rangle$. After scaling, we combine the spectra piecewise at the join wavelengths (no additional cross-fade). The resulting template covers 90 to 2100 nm on a uniform grid.
Appendix D Example cutouts of quasars
@@
Figure 15: Euclid 1D spectra and imaging cutouts for a random sample of 10 low-redshift ($z<0.5$) sources. Each panel displays the 1D spectrum (top) and five imaging cutouts with $10^{\prime\prime}$ sizes (bottom) in the $I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ bands, as well as a VIS-$Y_{\scriptscriptstyle\rm E}$ composite. The composite image is generated by mapping the VIS and $Y_{\scriptscriptstyle\rm E}$ fluxes into the blue and red channels, respectively, with their mean used for green, and the VIS band used to define overall luminosity in the L*a*b* colour space (International Organization for Standardization and International Commission on Illumination 2019) to enhance morphological detail. Major emission lines detected in the wavelength range [12 047, 18 734] Å are marked.